A conducting sphere has radius R_1 = 0.5m Calcula
Answer #1

(a) Capacitance is given by

C=4*pi*epsilon*radius=5.56*10^(-11) F

(b) if two sphere

let capacitance be C'

AV 1

if k=1485

capacitance=C'=kC=1.65*10^(-7) F